OpenStudy (anonymous):

Mrs. Guzman bought3table lamps for$129.90. If each lamp cost $40.00 before tax was added, what tax rate did she pay on the 3 lamps?

OpenStudy (anonymous):

Forget about the tax part for now. So Mrs. Guzman bought 3 table lamps at $40 each. How much should she have paid?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

129.90 = 3(40) + r (3*40) r = 0.0825 or 8.25&

OpenStudy (anonymous):

*%*

OpenStudy (anonymous):

$40*3 = $120.00 The total was $129.90. So the difference is 9.90. That's the tax that she paid. To find the rate, do this: If 120.00 = 100% Then 29.90 = ? (100%*9.90)/120.00 = 8.25%. Therefore, the Tax rate is 8.25%

OpenStudy (anonymous):

thanks ;)

OpenStudy (anonymous):

each lamp costs $40, the cost of the three lamps together without the tax included would be $120. so from there you want to subtract 129.9-120 to get 9.9. That number (the 9.9) is all of the money you paid simply for tax. So if you divide that by three (because there are three lamps) you would get 3.3 (meaning you pay $3.30 per each lamp) from there you simply divide 3.3 by 40 to get 8.25%

OpenStudy (anonymous):

\[3 (40 x+40)=\frac{12990}{100} \]If you solve the above for x you will obtain the same tax rate as @mangorox.